A car moving at 24 m/s skids to a stop in a distance of 30. meters. How long (in seconds) did it take for the car to come to a stop?
1
Expert's answer
2020-09-01T10:37:27-0400

"S=\frac{v^2-v_0^2}{2a}\to a=\frac{v_0^2}{2S}"


"v=v_0-at\to t=\frac{v_0}{a}=\frac{v_0\cdot2S}{v_0^2}=\frac{2S}{v_0}=\frac{2\cdot30}{24}=2.5(s)"